OpenStudy (anonymous):

Choose the polynomial that is written in standard form. −3x5y2 + 4x3y + 10x2 −8xy2 + 4x4y2 + 3x3 x4y2 + 4x3y5 + 10x4 x6y2 + 4x3y8 + 10x7 I just do get Algebra I but any other math i can do lol help me please.

OpenStudy (anonymous):

standard form means highest degree to lowest to get the degree of each term, add the exponents for example the degree of the term \(x^4y^2\) is \(4+2=6\)

OpenStudy (luigi0210):

Standard form is: Ax+By=C

OpenStudy (ja1):

add all the exponents in their terms and then see if they are in declining order.

OpenStudy (anonymous):

i know the standard from but lost after that

OpenStudy (ja1):

For example: \[−3x^5y^{2} + 4x^3y + 10x^2\]

OpenStudy (ja1):

Add the exponents up

OpenStudy (anonymous):

so I will had to add just like is right?????

OpenStudy (ja1):

Like @satellite73 said

OpenStudy (ja1):

But only add the exponents in their relevant terms

OpenStudy (anonymous):

\[ \overbrace{−8xy^2}^3 +\overbrace{ 4x^4y^2}^6 +\overbrace{ 3x^3}^3 \]

OpenStudy (ja1):

^

OpenStudy (anonymous):

not in standard from because it does not go from highest to lowest

OpenStudy (ja1):

Remember any lone variable (y,x) is the same as x^1, y^1

OpenStudy (ja1):

One way I do it is to look at the first two terms, if the first one is smaller immediately move one.

OpenStudy (anonymous):

\[\overbrace{ x^4y^2}^6 + \overbrace{4x^3y^5}^8 +\overbrace{ 10x^4}^4 \]also no

OpenStudy (ja1):

on*

OpenStudy (anonymous):

so you mean x4 the 4 is bigger then the x. i don't get this

OpenStudy (anonymous):

add up the exponents in each term

OpenStudy (anonymous):

forget about the variable, just add the exponents

OpenStudy (ja1):

Exactly, just add up the little numbers on top, and then see if they go down in size

OpenStudy (anonymous):

the ones on the bottom or the top

OpenStudy (ja1):

Top

OpenStudy (anonymous):

i got 18

OpenStudy (ja1):

\[x^{2} \]

OpenStudy (ja1):

? One which one?

OpenStudy (anonymous):

18 for the top

OpenStudy (anonymous):

lets go slow start with \[ −3x^5y^2 + 4x^3y + 10x^2 \] which has 3 terms

OpenStudy (ja1):

Every part seperated by the + or - sign is a term (-3x^5y^2) is a term

OpenStudy (anonymous):

one term is \(-3x^5y^2\) and the degree of that term you get by adding the exponents ignore everything else the degree of \(-3x^5y^2\) is \(7\) because \(2+5=7\)

OpenStudy (anonymous):

how about the degree of \(4x^3y\) ?

OpenStudy (anonymous):

4+2=6 right

OpenStudy (ja1):

Not the big number, only the exponents :)

OpenStudy (anonymous):

I am so lost :(

OpenStudy (anonymous):

how about this example: the degree of \(x^3y^7\) is \(3+7=10\) is that ok?

OpenStudy (anonymous):

how did you do that

OpenStudy (anonymous):

i added the exponent of \(x^3\), which is 3, to the exponent of \(y^7\) which is 7 when you add 3 and 7 you get 10

OpenStudy (anonymous):

ok with that?

OpenStudy (anonymous):

x2 and y7 how do you add them to 3 and 7 to = 10

OpenStudy (ja1):

I thought that was a two for a second <_< Anyways felicia, you see an exponent is the small little number on top of the variables (x, y are the variables) and the big number at the beginning is irrelevant.

OpenStudy (anonymous):

if you had \(x^2y^7\) then the degree would not be 10, it would be 9, because \(2+7=9\)

OpenStudy (anonymous):

When i did i got 9

OpenStudy (anonymous):

But i do know how you got 10

OpenStudy (anonymous):

ok good, now how about \[17x^3y^7\] what do you think the degree of that is?

OpenStudy (anonymous):

i add the top or the bottom???

OpenStudy (anonymous):

just the exponents, that is the point ignore the 17 entirely and just add the exponents

OpenStudy (anonymous):

Will you get 9????

OpenStudy (anonymous):

not when you add 3 and 7, no you still get 10

OpenStudy (anonymous):

how about \(-5x^4y^2\) ?

OpenStudy (anonymous):

i forget the 2 is a 3 sorry

OpenStudy (anonymous):

lol no problem

OpenStudy (anonymous):

what about \(-5x^4y^2\) what do you think the degree of that one is?

OpenStudy (anonymous):

so will 4 be 5 now???

OpenStudy (anonymous):

exponents of \(-5x^4y^2\) are \(4\) and \(2\) add them up

OpenStudy (anonymous):

4+2=6

OpenStudy (anonymous):

ok got that one how about \[−3x^5y^2\]

OpenStudy (anonymous):

is the 2 a 3 now???

OpenStudy (anonymous):

ok i see you are confused a bit there is no such thing as "the 2 is a 3" read what you see as the exponents, and add them

OpenStudy (anonymous):

5+2=7

OpenStudy (anonymous):

don't change the exponents, add them right!

OpenStudy (anonymous):

now what about \( 10x^2\) ? there is only one exponent here, so nothing to add

OpenStudy (anonymous):

so the degree of \(10x^2\) is just \(2\)

OpenStudy (ja1):

Ok try this, ignore everything except for the small numbers on top.

OpenStudy (anonymous):

why is just 2

OpenStudy (anonymous):

because that is the only exponent, there is no other

OpenStudy (anonymous):

the degree of \(x^4\) is \(4\) and the degree of \(15x^6\) is \(6\)

OpenStudy (anonymous):

ok get it

OpenStudy (anonymous):

because you can't add nothing to it right

OpenStudy (anonymous):

right

OpenStudy (anonymous):

ok one more what do you think the degree of \(4x^3y\) is? careful here

OpenStudy (anonymous):

will it be 2???

OpenStudy (anonymous):

a hint is \(4x^3y\) is the same as \(4x^3y^1\)

OpenStudy (anonymous):

what are the exponents?

OpenStudy (anonymous):

2+1=3

OpenStudy (ja1):

Remember any variable that is lone is the same as 1

OpenStudy (ja1):

Yupp

OpenStudy (anonymous):

YES

OpenStudy (anonymous):

now lets look at your first problem, which we have actually completed \[ −3x^5y^2 + 4x^3y + 10x^2 \]

OpenStudy (ja1):

Breakthrough :D

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

you have already said what the degree of each term is there

OpenStudy (anonymous):

\[−3x^5y^2 + 4x^3y + 10x^2\] you already said the degree of \(-3x^5y^2\) is 7 right?

OpenStudy (anonymous):

and you just said the degree of \(4x^3y\) is \(4\)

OpenStudy (anonymous):

yes yes

OpenStudy (anonymous):

and a little earlier you said the degree of \(10x^2\) is \(2\)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

the the degrees are, in order, \(7,4,2\)

OpenStudy (anonymous):

and they are definitely ordered from largest to smallest right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then it is in standard form!

OpenStudy (anonymous):

ok so thats it????????

OpenStudy (ja1):

I admire your ways to get t hings done @satellite73 btw you know I can't PM you back right xD

OpenStudy (anonymous):

now you can

OpenStudy (anonymous):

@Felicia123 yes, that is it, unless there is another one in standard form too

OpenStudy (anonymous):

so the next one be like

OpenStudy (anonymous):

you might want to make sure you understand by finding the degrees of the terms of \[ x^6y^2 + 4x^3y^8 + 10x^7\]

OpenStudy (anonymous):

6+2=8 2+8=10 and just 7

OpenStudy (anonymous):

i guess you got it, except for one small detail for some reason you keep seeing the \(3\) as a \(2\)!